OpenStudy (anonymous):

A floodlight illuminates a tall vertical wall that is 6m from it. A man, 2m tall, walks towards the wall in front of the light. HELP! I need to find a function that gives the height of his shadow as a function of his distance from the light???

OpenStudy (amistre64):

application of derivatives :)

OpenStudy (amistre64):

how fast is he walking?

OpenStudy (anonymous):

nice art!

OpenStudy (amistre64):

i got a tattoo like that lol

OpenStudy (amistre64):

\[\frac{ds}{dt}=\frac{ds}{}\frac{}{dt}\] we need something to relate this to; most likely the speed of the walk \[\frac{ds}{dt}=\frac{ds}{dw}\frac{dw}{dt}\] or some combo if chains

OpenStudy (amistre64):

6-d = x h = 2 s^2 = (6-d)^2 + 2^2

OpenStudy (amistre64):

im thinking of this as how fast the shadow is moving; wrong question altogether

OpenStudy (amistre64):

the shadow itself as a function

OpenStudy (amistre64):

you have similar triangles; 6 is to 6-(distance from light) as 2 is to (height of shadow)

OpenStudy (anonymous):

ya, I'm confused. There is no rate given?!?! That literally is what the question says.....I'm doing a Calculus course by correspondence....frustrating to say the least! The next question says "what is his shadow height when he is 1m from the light?"...and i| know the answer is 12?!?!? does that help??

OpenStudy (anonymous):

but 6 is a horizontal measure as well as x...I was thinking hte same thing but having a mental block...can't make up the triangle, as I dont' know the height of the light?

OpenStudy (amistre64):

the ratio would be: \[\frac{6}{2}=\frac{6-d}{sh}\]

OpenStudy (amistre64):

the light is at ground level; its a flood light

OpenStudy (amistre64):

my 6-d is wrong; it should simply be d

OpenStudy (amistre64):

\[\frac{6}{2}=\frac{d}{sh}\frac{\text{ distance from light}}{\text{ height of shadow}}\] the ratios are good since we are ddealing with similar triangles right?

OpenStudy (amistre64):

height of his shadow as a function of his distance from the light got to revise it still \[\frac{6}{sh}=\frac{d}{2}\] thats better

OpenStudy (amistre64):

shadow height = 12/d

OpenStudy (amistre64):

does that make sense to you?

OpenStudy (amistre64):

similar triangles produce a scalar effect; so we ratio the measurements

OpenStudy (anonymous):

sort of....it makes sense because I understand that I needed to set up ratios, and I know the answer when d=1 is 12, but why match the 6 with the s and the d with his height? how will I know for future questions which ones match up?

OpenStudy (amistre64):

the wall never moves; so its always at a distance of 6; the shadow is always on the wall, so its gots to be part of that specific triangle the height of the man never changes; so it has to be part of the triangle that moves as the man moves

OpenStudy (amistre64):

man height (2) is to shadow height as distance is to wall distance (6) \[\frac{2}{d}=\frac{sh}{6}\] solve for sh: sh = 12/d

OpenStudy (anonymous):

ok, got it, thanks! next questions is to find a function that gives the rate of change of the shadow, so, I guess differentiate it,,,,,correct?

OpenStudy (amistre64):

rate of change with respect to what? time or distance?

OpenStudy (anonymous):

distance

OpenStudy (amistre64):

might as well do it both ways :)

OpenStudy (amistre64):

wrt distance we get \[\frac{d}{dd}(sh=12/d)\implies\frac{d\ sh}{dd}=\frac{-12}{d^2}*\frac{dd}{dd}\]

OpenStudy (amistre64):

wrt time its pretty much the same: \[\frac{d}{dt}(sh=12/d)\implies\frac{d\ sh}{dt}=\frac{-12}{d^2}*\frac{dd}{dt}\] notice the denominator "dd" is now "dt"

OpenStudy (amistre64):

dd/dt is the speed at which the guy walks

OpenStudy (anonymous):

what's dd/dd in above answer? I got -12/d^2 but didn't write dd/dd...is that wrong?

OpenStudy (amistre64):

its a part of derivatives that they never teach you, hoping to confuse you later on ... recall that the chain rule always pops out a derived bit. for example: \[\frac{d}{dx}(5y^3)=15y^2*\frac{dy}{dx}\text{ ;the chain rule poped out a derived bit of dy/dx}\] \[\frac{d}{dx}(6x^2)=12x*\frac{dx}{dx}\text{ ; since dx/dx = 1 they tend to omit it}\]

OpenStudy (amistre64):

dd/dd is just the derived bit of distance with respect to distance

OpenStudy (amistre64):

it would do you good to keep all your derived bit intact until the end so that you can sort thru them afterwards. make sense?

OpenStudy (anonymous):

yup. thanks so much! probably be asking for more help shortly....we'll see! lol!